# Source Code Summarization in the Era of Large Language Models

### v1

@@ IV-A RQ1: What evaluation methods are suitable for assessing the quality of summaries generated by LLMs?
2) Experimental Results. Table V presents the scores reported by the GPT-4 evaluation method for summaries generated by four LLMs using five prompting techniques. Observe that when the base model is CodeLlama-Instruct, few-shot prompting consistently performs best on all three datasets. When the base model is StarChat-$\beta$, chain-of-thought prompting performs best on all the Java and C datasets, while expert prompting excels on the Python dataset. When selecting GPT-3.5 as the base model, the simplest zero-shot prompting surprisingly achieves the highest scores on the Java and C datasets, and is only slightly worse than few-shot prompting on the Python dataset. When using GPT-4 as the base model, chain-of-thought prompting overall performs best.
@@ IV-C RQ3: How do different model settings affect LLMs’ code summarization performance?
Top_p: In each round of token generation, LLMs sort tokens by probability from high to low and keep tokens whose probability adds up to (no more than) top_p. For example, $top\_p=0.1$ means only the tokens comprising the top 10% probability mass are considered. The larger the top_p is, the more tokens are sampled. Thus tokens with low probabilities have a greater chance of being selected, so the summary generated by LLMs is more random.
@@
TABLE VI: Influence of different model settings.We bold the scores of the best setting combinations on each dataset.
@@
2) Experimental Results. TABLE VI shows the scores evaluated by the GPT-4 evaluation method for the summaries generated by LLMs under different top_p and temperature settings. It is observed that the impact of top_p and temperature on the quality of LLM-generated summaries is specific to the base LLM and programming language. For example, when top_p=0.5, as temperature increases, the quality of GPT-4-generated summaries for Python code snippets increases, while those for C code snippets decrease. Another example is that when top_p=0.5, as the temperature rises, the quality of GPT-4-generated Java comments first increases and then decreases, whereas CodeLlama-Instruct is exactly the opposite, first decreases and then increases. Regarding the difference in influence between top_p and temperature, it is observed that in most cases the influence of the two parameters is similar. For example, for C code snippets, when one parameter (top_p or temperature) is fixed, as the other parameter (temperature or top_p) grows, the quality of GPT-3.5-generated summaries first decreases and then increases.
@@
TABLE VII: Effectiveness of LLMs in summarizing code snippets written in different types of programming languages. CodeLlama-I: CodeLlama-Instruct.
Model OOP PP SP FP LP
@@
StarChat-$\beta$ 2.50 2.68 2.97 2.37 2.79 2.73 2.67 2.68 2.88 2.34
GPT-3.5 3.73 3.56 4.14 3.97 3.64 3.99 3.53 3.57 3.44 3.42
GPT-4 4.38 4.18 4.36 4.16 4.37 4.31 4.29 4.23 4.22 4.05
2) Experimental Results. Table VII shows the performance evaluated by the GPT-4 evaluation method for the four LLMs on five types of programming languages. It is observed that for OOP (i.e., Java), GPT-4 performs best, followed by CodeLlama-Instruct, GPT-3.5, and StarChat-$\beta$. For PP, GPT-4 performs best on both C and Go, while StarChat-$\beta$ performs worst on both. The smallest LLM CodeLlama-Instruct outperforms GPT-3.5 on C (3.91 vs. 3.56), but vice versa on Go (3.86 vs. 4.14). Additionally, except for CodeLlama-Instruct, which performs slightly worse on Go than on C (3.86 vs. 3.91), the other three LLMs perform better on Go than on C. For SP, GPT-4 consistently performs best on all four languages. Surprisingly, CodeLlama-Instruct outperforms GPT-3.5 on both Ruby and JavaScript. All four LLMs perform better on PHP than on Python. For FP, the performance of two specialized code LLMs (i.e., CodeLlama-Instruct and StarChat-$\beta$) is better on Haskell than on Erlang, while the opposite is true for the two general-purpose LLMs (i.e., GPT-3.5 and GPT-4). For LP, GPT-4 still performs best, followed by GPT-3.5, CodeLlama-Instruct, and StarChat-$\beta$. Across all five types of languages, the four LLMs consistently perform the worst on LP, which indicates that summarizing logic programming language code is the most challenging. One possible reason is that fewer Prolog datasets are available for training these LLMs compared to other programming languages. The scale of the Prolog dataset we collected can support this reason.
@@ IV-E RQ5: How do LLMs perform on different categories of summaries?
TABLE VIII: Statistics of six sub-datasets divided from the CSN-Java test dataset according to comment intention
@@
     Others      37      $<0.01$
We employ the comment classifier COIN provided by Mu et al. [55] to classify the CSN-Java dataset according to the comment intention type. The test dataset is divided into six sub-datasets, as shown in Table VIII. To facilitate comparison between different categories, we randomly select 180 samples from each sub-dataset. As in RQ4, we uniformly employ few-shot prompting to adapt LLMs. For each sub-dataset with different intention types, the few-shot example is of the same intention type from the training dataset.
TABLE IX: Effectiveness of LLMs in generating different categories of summaries
@@
2) Experimental Results. Table IX presents the results evaluated by the GPT-4 evaluation method for the four LLMs in generating five categories of summaries. Observe that CodeLlama-Instruct performs worse in generating How-it-is-done summaries than generating the other four categories of summaries. StarChat-$\beta$ gets the lowest score of 2.52 in generating How-to-use summaries. Both GPT-3.5 and GPT-4 are not as good at generating Property summaries compared to generating other categories of summaries. Surprisingly, the smallest LLM CodeLlama-Instruct slightly outperforms the advanced GPT-4 in generating Why (4.29 vs. 4.28) and Property (4.19 vs. 4.06) summaries. Additionally, compared with GPT-3.5, CodeLlama-Instruct achieves higher scores in generating What, Why, and Property summaries. Certainly, it is undeniable that the reason for this phenomenon is that the optimal prompting technique for GPT-3.5 and GPT-4 is not few-shot prompting. This phenomenon is also exciting because it implies that most ordinary developers or teams who lack sufficient resources (e.g., GPUs) have the opportunity to utilize open-source and small-scale LLMs to achieve code summarization capabilities close to (or even surpass) those of commercial gigantic LLMs.
@@ Acknowledgment
The authors would like to thank the anonymous reviewers for their insightful comments. This work is supported partially by the National Natural Science Foundation of China (61932012, 62372228), and the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore.

### v2

@@ IV-A RQ1: What evaluation methods are suitable for assessing the quality of summaries generated by LLMs?
2) Experimental Results. Table IV presents the scores reported by the GPT-4 evaluation method for summaries generated by four LLMs using five prompting techniques. Observe that when the base model is CodeLlama-Instruct, few-shot prompting consistently performs best on all three datasets. When the base model is StarChat-$\beta$, chain-of-thought prompting performs best on all the Java and C datasets, while expert prompting excels on the Python dataset. When selecting GPT-3.5 as the base model, the simplest zero-shot prompting surprisingly achieves the highest scores on the Java and C datasets, and is only slightly worse than few-shot prompting on the Python dataset. When using GPT-4 as the base model, chain-of-thought prompting overall performs best.
@@ IV-C RQ3: How do different model settings affect LLMs’ code summarization performance?
Top_p: In each round of token generation, LLMs sort tokens by probability from high to low and keep tokens whose probability adds up to (no more than) top_p. For example, top_p=0.1top\_p=0.1 means only the tokens comprising the top 10% probability mass are considered. The larger the top_p is, the more tokens are sampled. Thus tokens with low probabilities have a greater chance of being selected, so the summary generated by LLMs is more random.
@@
TABLE V: Influence of different model settings. We bold the scores of the best setting combinations on each dataset.
@@
2) Experimental Results. TABLE V shows the scores evaluated by the GPT-4 evaluation method for the summaries generated by LLMs under different top_p and temperature settings. It is observed that the impact of top_p and temperature on the quality of LLM-generated summaries is specific to the base LLM and programming language. For example, when top_p=0.5, as temperature increases, the quality of GPT-4-generated summaries for Python code snippets increases, while those for C code snippets decrease. Another example is that when top_p=0.5, as the temperature rises, the quality of GPT-4-generated Java comments first increases and then decreases, whereas CodeLlama-Instruct is exactly the opposite, first decreases and then increases. Regarding the difference in influence between top_p and temperature, it is observed that in most cases the influence of the two parameters is similar. For example, for C code snippets, when one parameter (top_p or temperature) is fixed, as the other parameter (temperature or top_p) grows, the quality of GPT-3.5-generated summaries first decreases and then increases.
@@
TABLE VI: Effectiveness of LLMs in summarizing code snippets written in different types of programming languages. CodeLlama-I: CodeLlama-Instruct.
Model OOP $P$ SP FP LP
@@
StarChat-$\beta$ 2.60 2.68 2.97 2.37 2.79 2.73 2.67 2.68 2.88 2.34
GPT-3.5 3.56 3.56 4.14 3.97 3.64 3.99 3.53 3.57 3.44 3.42
GPT-4 4.66 4.18 4.36 4.16 4.37 4.31 4.29 4.23 4.22 4.05
2) Experimental Results. Table VI shows the performance evaluated by the GPT-4 evaluation method for the four LLMs on five types of programming languages. It is observed that for OOP (i.e., Java), GPT-4 performs best, followed by CodeLlama-Instruct, GPT-3.5, and StarChat-$\beta$. For PP, GPT-4 performs best on both C and Go, while StarChat-$\beta$ performs worst on both. The smallest LLM CodeLlama-Instruct outperforms GPT-3.5 on C (3.91 vs. 3.56), but vice versa on Go (3.86 vs. 4.14). Additionally, except for CodeLlama-Instruct, which performs slightly worse on Go than on C (3.86 vs. 3.91), the other three LLMs perform better on Go than on C. For SP, GPT-4 consistently performs best on all four languages. Surprisingly, CodeLlama-Instruct outperforms GPT-3.5 on both Ruby and JavaScript. All four LLMs perform better on PHP than on Python. For FP, the performance of two specialized code LLMs (i.e., CodeLlama-Instruct and StarChat-$\beta$) is better on Haskell than on Erlang, while the opposite is true for the two general-purpose LLMs (i.e., GPT-3.5 and GPT-4). For LP, GPT-4 still performs best, followed by GPT-3.5, CodeLlama-Instruct, and StarChat-$\beta$. Across all five types of languages, the four LLMs consistently perform the worst on LP, which indicates that summarizing logic programming language code is the most challenging. One possible reason is that fewer Prolog datasets are available for training these LLMs compared to other programming languages. The scale of the Prolog dataset we collected can support this reason.
@@ IV-E RQ5: How do LLMs perform on different categories of summaries?
TABLE VII: Statistics of six sub-datasets divided from the CSN-Java test dataset according to comment intention
@@
   Others    37    <0.01<0.01
We employ the comment classifier COIN provided by Mu et al. [44] to classify the CSN-Java dataset according to the comment intention type. The test dataset is divided into six sub-datasets, as shown in Table VII. To facilitate comparison between different categories, we randomly select 180 samples from each sub-dataset. As in RQ4, we uniformly employ few-shot prompting to adapt LLMs. For each sub-dataset with different intention types, the few-shot example is of the same intention type from the training dataset.
TABLE VIII: Effectiveness of LLMs in generating different categories of summaries
@@
2) Experimental Results. Table VIII presents the results evaluated by the GPT-4 evaluation method for the four LLMs in generating five categories of summaries. Observe that CodeLlama-Instruct performs worse in generating How-it-is-done summaries than generating the other four categories of summaries. StarChat-$\beta$ gets the lowest score of 2.52 in generating How-to-use summaries. Both GPT-3.5 and GPT-4 are not as good at generating Property summaries compared to generating other categories of summaries. Surprisingly, the smallest LLM CodeLlama-Instruct slightly outperforms the advanced GPT-4 in generating Why (4.29 vs. 4.28) and Property (4.19 vs. 4.06) summaries. Additionally, compared with GPT-3.5, CodeLlama-Instruct achieves higher scores in generating What, Why, and Property summaries. Certainly, it is undeniable that the reason for this phenomenon is that the optimal prompting technique for GPT-3.5 and GPT-4 is not few-shot prompting. This phenomenon is also exciting because it implies that most ordinary developers or teams who lack sufficient resources (e.g., GPUs) have the opportunity to utilize open-source and small-scale LLMs to achieve code summarization capabilities close to (or even surpass) those of commercial gigantic LLMs.
@@ Acknowledgment
The authors would like to thank the anonymous reviewers for their insightful comments. This work is supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), and the National Natural Science Foundation of China (61932012, 62372228). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and Cyber Security Agency of Singapore. Chunrong Fang is the corresponding author.